# Jacobian Descent For Multi-Objective Optimization

### v1

@@ Appendix C Experimental settings
For all of our experiments, we used PyTorch [55]. We have developed an open-source library [^4] Available at https://github.com/TorchJD/torchjd on top of it to enable Jacobian descent easily. This library is designed to be reusable for many other use cases than the experiments presented in our work. To separate them from the library, the experiments have been conducted with a different code repository [^5] Available at https://github.com/ValerianRey/jde mainly using PyTorch and our library.
@@ C.1 Learning rate selection
The learning rate has a very important impact on the speed of optimization. To make the comparisons as fair as possible, we always show the results corresponding to the best learning rate. We have selected the area under the loss curve as the criterion to compare the learning rates. This choice is arbitrary but seems to work well in practice: lower area under the loss curve means that the optimization is fast (quick loss decrease) and stable (few bumps in the loss curve). Concretely, for each random rerun and for each aggregator, we first try 22 learning rates from $10^{-5}$ to $10^{2}$, increasing by a factor $10^{\frac{1}{3}}$ every time. The two best learning rates from this range then define a refined range of plausible good learning rates, going from the smallest of those two multiplied by $10^{-\frac{1}{3}}$ to the largest of those two multiplied by $10^{\frac{1}{3}}$. This margin makes it unlikely for the best learning rate to lie out of the refined range. After this, 50 learning rates from the refined range are tried. These learning rates are evenly spaced in the exponent domain. The one with the best area under the loss curve is then selected and presented in the plots. Note that for simplicity, we have always used a constant learning rate, i.e. no learning rate scheduler was used.
We think that this approach has the advantage of being quite simple and very precise, thus giving trustworthy results. However, it requires a total of 72 trainings for each aggregator, each random rerun and each dataset. For this reason, we have opted for working on small subsets of the original datasets.
@@ C.2 Random reruns and standard error of the mean
To get an idea of confidence in our results, every experiment is performed 8 times on a different seed and on a different subset, of size 1024, of the training dataset. The seed used for run $i\in[8]$ is always simply set to $i$. Because each random rerun includes the full learning rate selection method described in Appendix C.1, it is sensible to consider the 8 sets of results as i.i.d. For each point of both the loss curves and the cosine similarity curves, we thus compute the estimated standard error of the mean with the usual formula $\frac{1}{\sqrt{8}}\sqrt{\frac{\sum_{i=1}^{8}(v_{i}-\bar{v})^{2}}{8-1}}$, where $v_{i}$ is the value of a point of the curve for random rerun $i$, and $\bar{v}$ is the average value of this point over the 8 runs.
@@ C.3 Model architectures
In all experiments, the models are simple convolutional neural networks. All convolutions always have a stride of 1x1, a kernel size of 3x3, a learnable bias and no padding. All linear layers always have a learnable bias. The activation function is the exponential linear unit [56]. The full architectures are given in Tables 234 and 5. Note that these architectures have been fixed arbitrarily, i.e. they were not optimized through some hyper-parameter selection. The weights of the model have been initialized with the default initialization scheme of PyTorch.
@@
Conv2d (3 input channels, 16 output channels, 1 groups), ELU
@@
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
@@
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
@@
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
@@
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
@@
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
@@
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
@@
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
@@
MaxPool2d (stride of 2x2, kernel size of 2x2), ELU
@@
MaxPool2d (stride of 3x3, kernel size of 3x3), ELU, Flatten
@@ C.4 Optimizer
For all experiments except those described in Appendix D.3, we always use the basic SGD optimizer of PyTorch, without any regularization or momentum. Note that here, SGD refers to the PyTorch optimizer that updates the parameters of the model in the opposite direction of the gradient, which, in our case, is replaced by the aggregation of the Jacobian matrix. In the rest of this paper, SGD refers to the whole stochastic gradient descent algorithm. In the experiments of Appendix D.3, we use Adam instead, to study its interactions with JD.
@@ C.6 Preprocessing
The inputs are always normalized per-channel based on the mean and standard deviation computed on the entire training split of the dataset.
@@ C.7 Iterations and computational budget
The numbers of epochs and the corresponding numbers of iterations for all datasets are provided in Table 6, along with the required number of NVIDIA L4 GPU-hours, to run all 72 learning rates for the 11 aggregators, on a single seed. The total computational budget to run the main experiments on 8 seeds was thus around 760 GPU-hours. Additionally, we have used a total of about 100 GPU-hours for the experiments varying the batch size and using Adam, and about 200 more GPU-hours were used for early investigations.
Table 6: Numbers of epochs, iterations and GPU-hours for each dataset
@@ D.1 All datasets and all aggregators
It should be noted that some of these aggregators were not developed as general-purpose aggregators, but mostly for the use case of multi-task learning, with one gradient per task. Our experiments present a more challenging setting than multi-task learning optimization, because conflict between rows of the Jacobian is typically higher. Besides, for some aggregators, e.g. $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{IMTL-G}}$, it was advised to make the aggregation of gradients w.r.t. an internal activation (such as the last shared representation), rather than w.r.t. the parameters of the model [17, 13]. To enable comparison, we instead always aggregated the Jacobian w.r.t. all parameters.
@@ D.2 Varying the batch size
Figure 9 shows the results on CIFAR-10 with $\mathcal{A}_{\text{UPGrad}}$ when varying the batch size from 4 to 64. Concretely, because we are using SSJD, this makes the number of rows of the sub-Jacobian aggregated at each step vary from 4 to 64. Recall that IWRM with SSJD and $\mathcal{A}_{\text{Mean}}$ is equivalent to ERM with SGD. We see that with a small batch size, in this setting, $\mathcal{A}_{\text{UPGrad}}$ becomes very similar to $\mathcal{A}_{\text{Mean}}$. This is not surprising, since with a batch size of 1, both would be equivalent. Conversely, a larger batch size increases the gap between $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$. Since the projections of $\mathcal{A}_{\text{UPGrad}}$ are onto the dual cone of more rows, each step becomes non-conflicting with respect to more of the original 1024 objectives, pushing even further the benefits of the non-conflicting property. In other words, increasing the batch size refines the dual cone, thereby improving the quality of the projections. It would be interesting to study more theoretically the impact of batch size in this setting.
@@ Appendix E Computation time
In this appendix, we compare the computation time of SGD with that of SSJD for all the aggregators that we experimented with. Since we used the same architecture for MNIST, Fashion-MNIST and Kuzushiji-MNIST, we only report the results for one of them. It is important to clarify all of the factors that affect this computation time. First, the batch size affects the number of rows in the Jacobian to aggregate. Increasing the batch size thus requires more GPU memory and the aggregation of a taller matrix. Then, some aggregators, e.g. $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{MGDA}}$, seem to greatly increase the run time. When the aggregation is the bottleneck, a faster implementation will be necessary to make them usable in practice. Lastly, the current implementation of our library for JD with PyTorch is still fairly inefficient in terms of memory management, which in turn limit how well the GPU can parallelize. Therefore, these results just give a rough indication of the current computation times, but they are in no way final.
@@
Appendix F Forward propagation of the Gramian through some typical layers
In this appendix, we study the forward propagation of the Gramian, as per (12), through some typical neural network layers. We start by noting that some computations can be saved when the function has a block structure.
Let ${\boldsymbol{g}}:{\mathbb{R}}^{n}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m}$ a function that maps an input ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$ and some parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$ to some output ${\boldsymbol{g}}({\boldsymbol{x}},{\boldsymbol{p}})$. Applying ${\boldsymbol{g}}$ independently in parallel yields a function ${\boldsymbol{f}}:{\mathbb{R}}^{n\times k}\times{\mathbb{R}}^{\ell}\to{\mathbb{R}}^{m\times k}$, where $k$ is the number of blocks. [^6] The block size $k$ could be, among others, the batch size or the number of elements of an activation. For an input $X\in{\mathbb{R}}^{n\times k}$ and a parameter ${\boldsymbol{p}}\in{\mathbb{R}}^{\ell}$,
${\boldsymbol{f}}(X,{\boldsymbol{p}})$ =[||$\ {\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{1},{\boldsymbol{p}})&\dots&{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{k},{\boldsymbol{p}})\\ |&&|\end{bmatrix}.$
In that case, forwarding the Gramian requires the matrices
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})$ $=\begin{bmatrix}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{1},{\boldsymbol{p}})&0&\cdots&0\\ 0&\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{2},{\boldsymbol{p}})&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{k},{\boldsymbol{p}})\end{bmatrix}$
$\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,{\boldsymbol{p}})$ $=\begin{bmatrix}J_{1}^{\vphantom{\top}}J_{1}^{\top}&J_{1}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{1}^{\vphantom{\top}}J_{k}^{\top}\\ J_{2}^{\vphantom{\top}}J_{1}^{\top}&J_{2}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{2}^{\vphantom{\top}}J_{k}^{\top}\\ \vdots&\vdots&\ddots&\vdots\\ J_{k}^{\vphantom{\top}}J_{1}^{\top}&J_{k}^{\vphantom{\top}}J_{2}^{\top}&\cdots&J_{k}^{\vphantom{\top}}J_{k}^{\top}\\ \end{bmatrix},$
where $J_{i}^{\vphantom{\top}}=\mathcal{J}\kern-2.0pt_{2}{\boldsymbol{g}}(X\cdot{\boldsymbol{e}}_{i},{\boldsymbol{p}})$.
Because of the block-diagonal structure of $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})$, for a given $G\in\mathcal{M}_{nk}$, the term $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})~{}\cdot~{}G\cdot~{}\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,{\boldsymbol{p}})^{\top}$ can be computed in $\mathcal{O}(k^{2}(mn^{2}~{}+~{}nm^{2}))$ operations rather than the naive $\mathcal{O}(k^{3}(mn^{2}~{}+~{}nm^{2}))$. Appendices F.1, F.2 and F.3 provide considerations about Gramian forwarding through some usual layers that have a particular block structure. For parameterized layers, $\mathcal{G}\kern-1.0pt_{2}f(X,{\boldsymbol{p}})$ can sometimes be precomputed efficiently, depending on the layer. More details are given in Appendix F.4 for a fully connected layer.
F.1 Activation functions
Activation functions generally consider each element of the activation independently. We thus have $m=n=1$, $\ell=0$ and $k$ equal to the number of elements of the activation multiplied by the batch size. The matrix $\mathcal{J}\kern-2.0pt_{1}f(X)$ is thus diagonal. The computational complexity of forwarding a Gramian through such a function is then $\mathcal{O}(k^{2})$. This applies to most usual activation functions, such as ReLU, ELU, sigmoid, and tanh.
F.2 Softmax layer
In the softmax layer, the group size $k$ is the batch size. Moreover, $m=n$ and $\ell=0$. The computational complexity of forwarding a Gramian is thus $\mathcal{O}(k^{2}m^{3})$.
F.3 BatchNorm layer
The batch normalization layer [57], in its simplest form (without the learnable affine transformation), normalizes its input on the batch dimension. Here $k$ is the input size, $\ell=0$, and $m=n$ is the batch size. In that case, the computational complexity of forwarding a Gramian is $\mathcal{O}(k^{2}m^{3})$.
F.4 Fully connected layer
In a fully connected layer, ${\boldsymbol{g}}:{\mathbb{R}}^{n}\times\mathbb{R}^{m\times n}\times{\mathbb{R}}^{m}\to{\mathbb{R}}^{m}$ is given by ${\boldsymbol{g}}({\boldsymbol{x}},W,{\boldsymbol{b}})=W{\boldsymbol{x}}+{\boldsymbol{b}}$ and ${\boldsymbol{f}}$ has a $k$-blockwise structure with $k$ the batch size. If $X\in{\mathbb{R}}^{n\times k}$ and $G\in{\mathbb{R}}^{nk\times nk}$, then forwarding those through the layer yields
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})^{\top}+\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})+\mathcal{G}\kern-1.0pt_{3}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$
with
$\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}W&0&\cdots&0\\ 0&W&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&W\end{bmatrix}$
$=I_{k}\otimes W,$
where $\otimes$ is the Kronecker product and $I_{k}$ the identity matrix of dimension $k$. In principle, $\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})\cdot G\cdot\mathcal{J}\kern-2.0pt_{1}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})^{\top}$ can be computed efficiently.
The Gramians with respect to the parameters are also rather simple. Let ${\boldsymbol{x}}_{i}=X{\boldsymbol{e}}_{i}$, then
$\mathcal{G}\kern-1.0pt_{2}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{1}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ {\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{2}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ \vdots&\vdots&\ddots&\vdots\\ {\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{1}I_{m}&{\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{2}I_{m}&\cdots&{\boldsymbol{x}}_{k}^{\top}{\boldsymbol{x}}_{k}I_{m}\\ \end{bmatrix}$
$=(X^{\top}X)\otimes I_{m}$
$\mathcal{G}\kern-1.0pt_{3}{\boldsymbol{f}}(X,W,{\boldsymbol{b}})$ $=\begin{bmatrix}I_{m}&I_{m}&\cdots&I_{m}\\ I_{m}&I_{m}&\cdots&I_{m}\\ \vdots&\vdots&\ddots&\vdots\\ I_{m}&I_{m}&\cdots&I_{m}\\ \end{bmatrix}$
$=1_{k}\otimes I_{m},$
with $1_{k}$ the all $1$ matrix of dimension $k\times k$.
Therefore, (12) rewrites as
$(I_{k}\otimes W)\cdot G\cdot(I_{k}\otimes W)+(X^{\top}X+1_{k})\otimes I_{m}.$
This illustrates that one should think of $G$ as the quadratic mapping $J\to JGJ^{\top}$, and similarly of $J$ as the linear mapping $G\to JGJ^{\top}$. Representing those mappings with arrays would be sub-optimal. Taking advantage of their structure (in this case, the Kronecker product of two matrices) would be beneficial.
It is worth noting that even though $(X^{\top}X+1_{k})\otimes I_{m}$ encompasses the dependencies of the Gramian to the parameters, it does not require any computation over the parameter dimension. If the number of parameters is large, this can provide a substantial improvement over the naive computation of the Gramian via the Jacobian.

### v2

@@ Appendix C Experimental settings
For all of our experiments, we used PyTorch (Paszke et al., 2019). We have developed an open-source library [^4] Available at https://github.com/TorchJD/torchjd on top of it to enable Jacobian descent easily. This library is designed to be reusable for many other use cases than the experiments presented in our work. To separate them from the library, the experiments have been conducted with a different code repository [^5] Available at anonymousGHlinkValerianReyjde mainly using PyTorch and our library.
@@ C.1 Learning rate selection
The learning rate has a very important impact on the speed of optimization. To make the comparisons as fair as possible, we always show the results corresponding to the best learning rate. We have selected the area under the loss curve as the criterion to compare the learning rates. This choice is arbitrary but seems to work well in practice: a lower area under the loss curve means the optimization is fast (quick loss decrease) and stable (few bumps in the loss curve). Concretely, for each random rerun and each aggregator, we first try 22 learning rates from $10^{-5}$ to $10^{2}$, increasing by a factor $10^{\frac{1}{3}}$ every time. The two best learning rates from this range then define a refined range of plausible good learning rates, going from the smallest of those two multiplied by $10^{-\frac{1}{3}}$ to the largest of those two multiplied by $10^{\frac{1}{3}}$. This margin makes it unlikely for the best learning rate to lie out of the refined range. After this, 50 learning rates from the refined range are tried. These learning rates are evenly spaced in the exponent domain. The one with the best area under the loss curve is then selected and presented in the plots. For simplicity, we have always used a constant learning rate, i.e. no learning rate scheduler was used.
This approach has the advantage of being simple and precise, thus giving trustworthy results. However, it requires 72 trainings for each aggregator, random rerun, and dataset, i.e. a total of $43776$ trainings for all of our experiments. For this reason, we have opted to work on small subsets of the original datasets.
@@ C.2 Random reruns and standard error of the mean
To get an idea of confidence in our results, every experiment is performed 8 times on a different seed and a different subset, of size 1024, of the training dataset. The seed used for run $i\in[8]$ is always simply set to $i$. Because each random rerun includes the full learning rate selection method described in Appendix C.1, it is sensible to consider the 8 sets of results as i.i.d. For each point of both the loss curves and the cosine similarity curves, we thus compute the estimated standard error of the mean with the usual formula $\frac{1}{\sqrt{8}}\sqrt{\frac{\sum_{i\in[8]}(v_{i}-\bar{v})^{2}}{8-1}}$, where $v_{i}$ is the value of a point of the curve for random rerun $i$, and $\bar{v}$ is the average value of this point over the 8 runs.
@@ C.3 Model architectures
In all experiments, the models are simple convolutional neural networks. All convolutions always have a stride of 1$\times$1, a kernel size of 3$\times$3, a learnable bias, and no padding. All linear layers always have a learnable bias. The activation function is the exponential linear unit (Clevert et al., 2015). The full architectures are given in Tables 234 and 5. Note that these architectures have been fixed arbitrarily, i.e. they were not optimized through some hyper-parameter selection. The weights of the model have been initialized with the default initialization scheme of PyTorch.
@@
Conv2d (3 input channels, 16 output channels, 1 group), ELU
@@
MaxPool2d (stride of 2$\times$2, kernel size of 2$\times$2), ELU
@@
MaxPool2d (stride of 3$\times$3, kernel size of 3$\times$3), ELU, Flatten
@@
MaxPool2d (stride of 2$\times$2, kernel size of 2$\times$ 2), ELU
@@
MaxPool2d (stride of 3$\times$3, kernel size of 3$\times$3), ELU, Flatten
@@
MaxPool2d (stride of 2$\times$2, kernel size of 2$\times$2), ELU
@@
MaxPool2d (stride of 2$\times$2, kernel size of 2$\times$2), ELU
@@
MaxPool2d (stride of 3$\times$3, kernel size of 3$\times$3), ELU, Flatten
@@
MaxPool2d (stride of 2$\times$2, kernel size of 2$\times$2), ELU
@@
MaxPool2d (stride of 3$\times$3, kernel size of 3$\times$3), ELU, Flatten
@@ C.4 Optimizer
For all experiments except those described in Appendix D.3, we always use the basic SGD optimizer of PyTorch, without any regularization or momentum. Here, SGD refers to the PyTorch optimizer that updates the parameters of the model in the opposite direction of the gradient, which, in our case, is replaced by the aggregation of the Jacobian matrix. In the rest of this paper, SGD refers to the whole stochastic gradient descent algorithm. In the experiments of Appendix D.3, we instead use Adam to study its interactions with JD.
@@ C.6 Preprocessing
The inputs are always normalized per channel based on the mean and standard deviation computed on the entire training split of the dataset.
@@ C.7 Iterations and computational budget
The numbers of epochs and the corresponding numbers of iterations for all datasets are provided in Table 6, along with the required number of NVIDIA L4 GPU-hours, to run all 72 learning rates for the 11 aggregators on a single seed. The total computational budget to run the main experiments on 8 seeds was thus around 760 GPU-hours. Additionally, we used a total of about 100 GPU hours for the experiments varying the batch size and using Adam, and about 200 more GPU hours were used for early investigations.
Table 6: Numbers of epochs, iterations, and GPU-hours for each dataset
@@ D.1 All datasets and all aggregators
It should be noted that some of these aggregators were not developed as general-purpose aggregators, but mainly for the use case of multi-task learning, with one gradient per task. Our experiments present a more challenging setting than multi-task learning optimization because conflict between rows of the Jacobian is typically higher. Besides, for some aggregators, e.g. $\mathcal{A}_{\text{GradDrop}}$ and $\mathcal{A}_{\text{IMTL-G}}$, it was advised to make the aggregation of gradients w.r.t. an internal activation (such as the last shared representation), rather than w.r.t. the parameters of the model (Chen et al., 2020; Liu et al., 2021b). To enable comparison, we instead always aggregated the Jacobian w.r.t. all parameters.
@@ D.2 Varying the batch size
Figure 9 shows the results on CIFAR-10 with $\mathcal{A}_{\text{UPGrad}}$ when varying the batch size from 4 to 64. Concretely, because we are using SSJD, this makes the number of rows of the sub-Jacobian aggregated at each step vary from 4 to 64. Recall that IWRM with SSJD and $\mathcal{A}_{\text{Mean}}$ is equivalent to ERM with SGD. We observe that with a small batch size, $\mathcal{A}_{\text{UPGrad}}$ becomes very similar to $\mathcal{A}_{\text{Mean}}$. This is not surprising since both would be equivalent with a batch size of 1. Conversely, a larger batch size increases the gap between $\mathcal{A}_{\text{UPGrad}}$ and $\mathcal{A}_{\text{Mean}}$. Since the projections of $\mathcal{A}_{\text{UPGrad}}$ are onto the dual cone of more rows, each step becomes non-conflicting with respect to more of the original 1024 objectives, pushing even further the benefits of the non-conflicting property. In other words, increasing the batch size refines the dual cone, thereby improving the quality of the projections. It would be interesting to theoretically analyze the impact of the batch size in this setting.
@@ Appendix E Computation time
In this appendix, we compare the computation time of SGD with that of SSJD for all the aggregators that we experimented with. Since we used the same architecture for MNIST, Fashion-MNIST and Kuzushiji-MNIST, we only report the results for one of them. Several factors affect this computation time. First, the batch size affects the number of rows in the Jacobian to aggregate. Increasing the batch size thus requires more GPU memory and the aggregation of a taller matrix. Then, some aggregators, e.g. $\mathcal{A}_{\text{Nash-MTL}}$ and $\mathcal{A}_{\text{MGDA}}$, seem to greatly increase the run time. When the aggregation is the bottleneck, a faster implementation will be necessary to make them usable in practice. Lastly, the current implementation of JD in our library is still fairly inefficient in terms of memory management, which in turn limits how well the GPU can parallelize. Therefore, these results just give a rough indication of the current computation times.